# Theories Without Models: Uncontrolled Idealizations in Particle Physics

### v1

@@ 5.2. Cutting Off Our Ignorance
The second problem that needs to be addressed in the construction of empirically relevant models in perturbative QFT is the presence of ultraviolet and infrared divergences in the individual terms of the power series in (5) which makes the calculation of the coefficients impossible. The elimination of these infinities is achieved by various methods of regularization in which a divergent integral $A_{n}$ is redefined as a function of a new parameter $\xi$ – the regulator – which must satisfy the two following pragmatic constraints: (i) finite values of the regulator must render the integral $A_{n}(\xi)$ finite and (ii) if the regulator is removed by taking its limit to infinity the retrieved result is the original divergent integral $A_{n}$. A standard regularization method is the Pauli-Villars regularization in which a cut-off limit $\Lambda$ is introduced for the ultraviolet and infrared domains of high and low energies respectively, beyond which the value of the integral is taken to be zero. [^12] Another popular method of regularization is dimensional regularization where, roughly speaking, the calculations are initially carried out in a d-dimensional spacetime, and the cut-off dependencies appear as we take the well-defined limit $d\rightarrow 4$. The Pauli-Villars method discussed here has the advantage of being physically more transparent as opposed to dimensional regularization which is considerably more abstract in nature. Some less popular methods of regularization are the lattice regularization, the zeta function regularization and the causal regularization. In the case of ultraviolet divergences, this is equivalent to the introduction of a mathematical idealization according to which
@@
where the values of $s_{0},t_{0}$ and $u_{0}$ are defined by the particular energy at which the coupling parameter $g_{R}$ was measured. At this stage, the dependence on the cut-off limit $\Lambda$ is dropped out, however, as one may notice the scattering amplitude now depends on the ratio between the energy scale $\mu$ at which the value of $g_{R}(\mu)$ is measured (captured by $s_{0},t_{0}$ and $u_{0}$) and the energy scale of the future scattering experiments by which the empirically relevant model will be tested (captured by $s,t$ and $u$). What is so special about the chosen energy scale $\mu$ however? And what if the renormalized parameter $g_{R}(\mu)$ was measured at a different scale $\mu^{\prime}$? One might think here that this is not much of an improvement since the dependence on the arbitrarily chosen value of $\Lambda$ has simply been shifted to a dependence on the energy scales $\mu$ at which we are able to conduct experiments and extract the value of the renormalized coupling parameter. The standard way to deal with this situation in the context of condensed matter physics is to think of Quantum Field Theory as a theoretical framework for providing Effective Field Theories that only describe physical phenomena at a particular range of energy scales (up to $\mu$) and not as a fundamental theory of physics. This idea leads us naturally to the framework of renormalization group (RG) theory (Wilson 1971; Fisher 1998). The renormalization group method provides the mathematical apparatus for a systematic investigation of the changes in the couplings of a theory with respect to the changes in the energy scales. Leaving the technical details aside, the main idea of this modern approach to renormalization is captured by the renormalization group flow equation
@@
    Safety. At a finite flow parameter value the renormalization group flow hits a point (also known as an attractor) at which the beta function becomes zero for all running couplings, $g_{i}(\mu)$. This is the so-called fixed point $*g*$ which basically implies that the theory becomes scale-invariant above a certain energy scale since all couplings converge to a fixed constant value. If a theory demonstrates this behaviour, it is said to be asymptotically safe.
@@
    Triviality The third possibility is the case in which the beta function is positive for at least one of the couplings in the theory and there is no non-trivial UV fixed point. This means that the running couplings $g_{i}(\mu)$ increase indefinitely as the energy scale increases. Theories without a non-trivial UV fixed point are considered trivial since their continuum limit is only well-defined if the values of all renormalized couplings, $g_{R}$, are zero. The physical interpretation of triviality is that the quantum corrections completely suppress the interactions in the absence of a finite cutoff: the physical charges are entirely ‘screened off’ if the ultraviolet cutoff is sent to infinity (Gies and Jaeckel 2004; Crowther and De Haro 2022). A particular problem associated with lack of a non-trivial continuum limit is where the renormalization group flow not only does not hit a fixed point but rather the beta function is found to blow up to infinity at a very large but finite energy level ($\mu_{Landau}$). Such a theory is said to encounter a Landau pole (Landau et al. 1954).
@@
Huggett (2002) presents a detailed discussion explaining how the renormalization group provides a map that helps us understand the infinities and the ‘maneuvering’ during the renormalization process, making the derived models consequences of the background theory. The basic idea is that one uses the renormalization group and a renormalization scale to generate a family of renormalized theories from a family of bare theories and then confirm that the family of bare theories describes well-defined physics in the continuum limit. The physics of every bare theory is captured by the physics of the corresponding renormalized theory, which is rendered finite by a fixed cut-off limit, and hence the physics of the limit of the bare theories is captured by the limit of the renormalized theories which is finite. The crucial step in the argument is to justify that we have well-defined continuum physics in the limit, and this is achieved, as (Huggett 2002, p.274) explains, by the topology of the parameter space. However, this explanation only holds for quantum field theories that are asymptotically safe such as QCD, where the renormalization group flow can be shown to hit a fix point. As Huggett puts it at the very end of his article: ‘…assuming that a theory is asymptotically safe, the problem is not terminal but can be fixed using renormalization to tune it to a suitable point. If, however, a theory is not asymptotically safe, then […] renormalization as described here will not be possible’ (ibid., p.275-6). [^16] Here it is worth mentioning the work of Franklin (2020) on the effectiveness of Effective Field Theories. Commenting on the paper of Butterfield (2014), Franklin notes that the appeal to the fixed point structure of the RG method in order to justify renormalization is ill suited. His claim is that RG is best seen as a mathematical framework that codifies rather than explains the renormalizability of theories in QFT, in that it allows us to mathematically establish whether or not a theory is renormalizable, but nonetheless, it does not provide a sufficient explanation of the physical reasons behind renormalization. Our focus here is on mathematical relationships rather than explanation and in Franklin’s terms, our point could be re-expressed as a statement that without a non-trivial UV fixed point the RG method cannot not successfully achieve its codifying role.
@@
Unfortunately, this rather attractive theoretical diagnostic picture is very difficult to fully implement in practice. For realistic QFTs there is no fully reliable way to explicitly calculate the beta functions to all orders. Take the example of QED. To render the introduced idealizations controllable, one needs to show that the theory is asymptotically safe by calculating the corresponding beta functions in realistic QED models, however, no rigorous proof for the asymptotic safety of QED has been achieved. Rather, what we have available are further mathematical idealization techniques for approximating the beta functions in some cases, which further suggest a possible behaviour of the theory in higher energies. Interestingly, the situation is then that we are attempting to justify the assumptions introduced by regularization and renormalization, by introducing further assumptions, creating a form of regress in which idealizations are justified by further idealizations. [^17] This is not to say that this is a vicious regress, however. Whether the justification of an approximation by means of a further approximation is epistemically warranted is an interesting open question and deserves to be studied on its own merit in future work.
@@
In any case, even ignoring the regress worry, the evidence gained from approximate methods for calculating the beta functions is not in fact promising for the case of QED. In particular, calculations at the one-loop level indicate the existence of a Landau pole (Hollowood 2013, p.37-41) and thus provide evidence against the asymptotic safety of QED . One might hope at this point that the Landau pole is an artifact of the perturbation calculation of the beta function and there is some evidence in this regard (Goeckeler et al. 1998). However, no indication of a non-trivial UV fixed point for QED has been found so far. Furthermore, evidence from non-perturbative lattice simulations (Goeckeler et al. 1998; Kim et al. 2001) and the exact renormalization group approach (Gies and Jaeckel 2004) supports the triviality of the theory with complete charge screening if the UV cutoff is sent to infinity.
@@
The standard response to the presumed lack of asymptotic safety in quantum field theories found in the physics literature is that such theories should not be understood as fundamental theories and a consistent UV completion can be obtained by appeal to more fundamental theories. However, it is not expected that electroweak theory can provide a UV completion of QED and, indeed, the triviality problem is expected to also occur in the Yukawa-Higgs sector. [^18] See for example Hambye and Riesselmann (1997) A full list of references is given in Eichhorn and Held (2018) reference [4]. Rather, the hope for a UV completion of QED relies upon appeal to physics beyond the standard model, such as in asymptotically safe quantum gravity scenarios (Christiansen and Eichhorn 2017; Eichhorn 2019). Whether or not such a response to the QED triviality problem ultimately proves successful is, however, tangential to the issue at hand: we are concerned with the justification for the idealizations involved in the theories and models considered in their own right, not exogenous justifications of the idealizations coming from other theories, cf. Crowther and De Haro (2022). Moreover, since this putative justification would need to come from theories we do not in fact have, the supposed resolution has the status of a promissory note. For the case of QED considered as a theory in its own right, the mathematical idealization in question is evidently uncontrolled (since no fixed point has been identified), and potentially also uncontrollable (since there is evidence that none in fact exists and the theory is trivial).
@@ 6. Implications
The analysis provided thus far has important implications for the philosophical understanding of theoretical models and the empirical adequacy of scientific theories. Let us recap the argument provided in Section 1. The first premise, P1, was that quantum field theory is a more general framework from which specific theories such as quantum electrodynamics and quantum chromodynamics are derived via particular action principles. The second premise, P2, was that a theory has theoretical models if its predicted empirical quantities are derivable either as deductive consequences of the theory or as a result of an approximate derivation involving controllable mathematical idealizations. The third premise, P3, was that the derivation of predictions from empirically successful theories in contemporary particle physics, such as quantum electrodynamics, necessarily involves appeal to uncontrollable idealizations. These three premises together lead us to the deeply unintuitive conclusion, C1, that theories in contemporary particle physics, such as quantum electrodynamics, do not have theoretical models.
The natural response for those seeking to reject the unintuitive conclusion is to challenge the controversial premise P3. However, as we saw in the previous section, P3 is, as it stands, remarkably resilient, even in the face of the sophisticated framework of renormalization group theory. Thus, it might seem that the only choice is to ‘bite the bullet’ and accept the conclusion of theories without models, despite its unintuitiveness. However, as we will argue now, this option is undesirable since C1 has further clearly unacceptable consequences for empirical adequacy. Consider the following extension of our argument:
@@
    Contemporary particle physics includes theories, such as quantum electrodynamics, without theoretical models.
@@
    A theory is empirically adequate only if it possesses theoretical models that agree with experimental outcomes.191919cf. (van Fraassen 1991, p.12): ‘To be empirically adequate is to have some model which can accommodate all the phenomena’.
@@
    The theory of quantum electrodynamics is not empirically adequate.
@@
Clearly, C2 is not just unintuitive but absurd and thus unacceptable. Quantum electrodynamics is variously understood as the most empirically successful theory humanity has yet constructed, so any argument that ends up with the conclusion that it is not empirical adequate must have gone wrong somewhere! Given that our robust defence of P3 is accepted and that P1 is also accepted as uncontroversial, the two remaining options to avoid absurdity are then to put pressure on P2 and P4. Let us consider each of these options in turn.
@@
If theoretical models are not required to be connected to their background theories via controlled idealizations, then one clearly avoids the absurd conclusion since one is able to understand the empirically relevant models of quantum electrodynamics as models of the theory. The cost to the coherence of our concepts is, however, catastrophic. This is because the expansion of the definition of a theoretical model to include links via uncontrollable mathematical idealizations, simpliciter, allows arbitrary mathematical connections between scientific theories and models. To put it differently, saying that the models of a theory include models which are the products of both controllable and uncontrollable idealizations allows any possible mathematical connection of any possible model with a background theory. Such a view would lead to a narrowly instrumentalist understanding of the empirical adequacy of theories which is solely defined in terms of the empirical success of the derived models, regardless of whether such models are constructed based upon rigorous and well justified scientific and mathematical processes or not. Consequently, under such view, any scientific theory in which various ad hoc mathematical modifications are introduced with the sole purpose of achieving agreement with data can be considered as empirically adequate insofar as the derived predictions agree with measurement outcomes. We do not, therefore, take the route via weakening P2 to be worth pursuing.
@@
In the context of the failure of Borel summability, the exogenous control approach is illustrated by the idea of appealing to non-perturbative ‘instanton’ corrections as a means to control the idealisation within the perturbative theory (Marino 2014). The key idea is to use properties of the non-perturbative solutions to understand the failure of Borel summability. In particular, we can differentiate two types of situation when there is no Borel summability: i) we are considering perturbative series around unstable minima; and ii) there are extra saddle points in the path integral. The second situation is understood to correspond to non-perturbative effects of the instanton type. In that context, one can consider ‘lateral Borel resummations’ (i.e. integrals along paths in the complex plane) and the ground state energy can be reconstructed from the Borel-resummed perturbative series corrected by Borel-resummed instanton solutions (Marino 2014, p. 37). The idea is, then, that realistic perturbative quantum field theories which are not Borel summable, admit a non-perturbative analysis with instanton corrections (or generalisations thereof). Formally, the relevant properties have only been shown to hold in toy models and thus the idealization in question is certainly not controlled as yet. However, there may be scope for experimental demonstration of the existence of instantons in the standard model (Amoroso et al. 2021) and this would go a long way towards provision of the kind of exogenous control that we take the idealizations of perturbative quantum field theory to require in order to anchor their empirical adequacy.
@@
There are also plausible possibilities for exogenous arguments that could provide control of the idealisations in the context of renormalization. These also come from a non-perturbative analysis but are exogenous to an even greater degree since they require us to include additional effects due to gravity. In particular, in the context of failure of asymptotic safety in quantum electrodynamics, scientists in the field of quantum gravity propose control of the relevant idealization by appeal to the existence of a fixed point in a unified theory of quantum gravity. This wider definition would be aligned to the general attitude that UV divergences of QFT point to quantum gravity as discussed by Crowther and De Haro (2022) and to the particular motivations for the asymptotic safety programme (Eichhorn 2019). [^20] See also Crowther and Linnemann (2019).
A related idea would be to try and offer a more relaxed definition of empirical adequacy based on a distinction between appropriate and inappropriate uncontrolled idealizations. Such an approach would require a robust framework for distinguishing between uncontrollable idealizations that are appropriate for the derivation of empirical predictions, such as the ones appearing in quantum electrodynamics, and other arbitrary and ad hoc idealizations to make models mathematically tractable. Although intuitively some idealizations seem plainly wrong, e.g. the introduction of an arbitrary number of free parameters to fit the data, the presented analysis of the status of idealizations in quantum electrodynamics indicates that there is a grey zone in which it is not clear whether a certain (uncontrolled) mathematical idealization is appropriate or not.

### v2

@@ 5.2. Cutting Off Our Ignorance
The second problem that needs to be addressed in the construction of empirically relevant models in perturbative QFT is the presence of ultraviolet and infrared divergences in the individual terms of the power series in (5) which makes the calculation of the coefficients impossible. The elimination of these infinities is achieved by various methods of regularization in which a divergent integral $A_{n}$ is redefined as a function of a new parameter $\xi$ – the regulator – which must satisfy the two following pragmatic constraints: (i) finite values of the regulator must render the integral $A_{n}(\xi)$ finite and (ii) if the regulator is removed by taking its limit to infinity the retrieved result is the original divergent integral $A_{n}$.
A standard regularization method is the Pauli-Villars regularization in which a cut-off limit $\Lambda$ is introduced for the ultraviolet and infrared domains of high and low energies respectively, beyond which the value of the integral is taken to be zero. [^16] Another popular method of regularization is dimensional regularization where, roughly speaking, the calculations are initially carried out in a d-dimensional spacetime, and the cut-off dependencies appear as we take the well-defined limit $d\rightarrow 4$. The Pauli-Villars method discussed here has the advantage of being physically more transparent as opposed to dimensional regularization which is considerably more abstract in nature. Some less popular methods of regularization are the lattice regularization, the zeta function regularization and the causal regularization. In the case of ultraviolet divergences, this is equivalent to the introduction of a mathematical idealization according to which
@@
An immediate implication of regularization is that the approximate value of the scattering amplitudes $\mathcal{M}$ now depends on the arbitrarily chosen value of $\Lambda$. However, given that there are no particular reasons for choosing a specific value for $\Lambda$ over a different one, say $\Lambda^{2}$, this should be worrying; different values of the arbitrarily chosen regulator will simply give different results for the scattering amplitudes, which is of course unacceptable. The scattering amplitudes provided by these models express the empirical content of our theory and thus they should not be a function of the theorists’ arbitrarily chosen value for $\Lambda$ [^19] It should be noted here that the choice of $\Lambda$ is not always completely arbitrary. For instance, when we want to abstract away from the interactions of a heavier, higher-energy particle, $\Lambda$ is chosen to be well below the mass of that particle and above the mass of the particles we are interested in studying. However, given that the range between the two masses is significantly large, the choice of the exact value for $\Lambda$ is still, more or less arbitrary. Butterfield and Bouatta (2015) also point out that in some cases the background theory hints towards a range of values for the cut-off limit that are plausible to take (p.448). An example of such a suggestion for the cut-off limit comes from the Lamb shift in QED, which suggests the electron’s Compton wavelength as a natural lower limit for distance $d$. Nonetheless, even in these cases no rigorous mathematical or physical justification can be given and thus the final choice of $\Lambda$ remains arbitrary.
The solution to this problem came from renormalization, cf. Fraser (2021). In this method the unwanted $\Lambda$-dependence of the integrals, and consequently of the scattering amplitudes, is eliminated by extracting the value of the coupling parameter $g$ from scattering experiments at a particular energy scale $\mu$. Once the value of the coupling parameter is obtained, the original ‘bare coupling’ $g_{0}$ is replaced by the new renormalized value $g_{R}(\mu)$ (sometimes called the ‘physical coupling parameter’) which is now a function of the energy scale $\mu$ at which the experiment was performed.
As a concrete example, consider the perturbation series for the scattering amplitude of a meson-meson scattering process up to the second order correction. Once the ultraviolet cut-off is introduced, the amplitude $\mathcal{M}$ becomes a finite ‘cut-off dependent’ quantity of the form
@@
where the values of $s_{0},t_{0}$ and $u_{0}$ are defined by the particular energy at which the coupling parameter $g_{R}$ was measured. At this stage, the dependence on the cut-off limit $\Lambda$ is dropped out, however, as one may notice the scattering amplitude now depends on the ratio between the energy scale $\mu$ at which the value of $g_{R}(\mu)$ is measured (captured by $s_{0},t_{0}$ and $u_{0}$) and the energy scale of the future scattering experiments by which the empirically relevant model will be tested (captured by $s,t$ and $u$).
What is so special about the chosen energy scale $\mu$ however? And what if the renormalized parameter $g_{R}(\mu)$ was measured at a different scale $\mu^{\prime}$? One might think here that this is not much of an improvement since the dependence on the arbitrarily chosen value of $\Lambda$ has simply been shifted to a dependence on the energy scales $\mu$ at which we are able to conduct experiments and extract the value of the renormalized coupling parameter. The standard way to deal with this situation in the context of condensed matter physics is to think of Quantum Field Theory as a theoretical framework for providing Effective Field Theories that only describe physical phenomena at a particular range of energy scales (up to $\mu$) and not as a fundamental theory of physics. This idea leads us naturally to the framework of renormalization group (RG) theory (Wilson 1971; Fisher 1998).
The renormalization group method provides the mathematical apparatus for a systematic investigation of the changes in the couplings of a theory with respect to the changes in the energy scales. Leaving the technical details aside, the main idea of this modern approach to renormalization is captured by the renormalization group flow equation
@@
    Safety. At a finite flow parameter value the renormalization group flow hits a point (also known as an attractor) at which the beta function becomes zero for all running couplings, $g_{i}(\mu)$. This is the so-called fixed point $g*$ which basically implies that the theory becomes scale-invariant above a certain energy scale since all couplings converge to a fixed constant value. If a theory demonstrates this behaviour, it is said to be asymptotically safe.
@@
    Triviality The third possibility is the case in which the beta function is positive for at least one of the couplings in the theory and there is no non-trivial UV fixed point. This means that the running couplings $g_{i}(\mu)$ increase indefinitely as the energy scale increases. Theories without a non-trivial UV fixed point are considered trivial since their continuum limit is only well-defined if the values of all renormalized couplings, $g_{R}$, are zero. The physical interpretation of triviality is that the quantum corrections completely suppress the interactions in the absence of a finite cut-off: the physical charges are entirely ‘screened off’ if the ultraviolet cut-off is sent to infinity (Gies and Jaeckel 2004; Crowther and De Haro 2022). A particular problem associated with lack of a non-trivial continuum limit is where the renormalization group flow not only does not hit a fixed point but rather the beta function is found to blow up to infinity at a very large but finite energy level ($\mu_{Landau}$). Such a theory is said to encounter a Landau pole (Landau et al. 1954).
@@
Huggett (2002) presents a detailed discussion explaining how the renormalization group provides a map that helps us understand the infinities and the ‘maneuvering’ during the renormalization process, making the derived models consequences of the background theory. The basic idea is that one uses the renormalization group and a renormalization scale to generate a family of renormalized theories from a family of bare theories and then confirm that the family of bare theories describes well-defined physics in the continuum limit. The physics of every bare theory is captured by the physics of the corresponding renormalized theory, which is rendered finite by a fixed cut-off limit, and hence the physics of the limit of the bare theories is captured by the limit of the renormalized theories which is finite. The crucial step in the argument is to justify that we have well-defined continuum physics in the limit, and this is achieved, as Huggett (2002, p.274) explains, by the topology of the parameter space. However, this explanation only holds for quantum field theories that are asymptotically safe such as QCD, where the renormalization group flow can be shown to hit a fix point. As Huggett puts it at the very end of his article: ‘…assuming that a theory is asymptotically safe, the problem is not terminal but can be fixed using renormalization to tune it to a suitable point. If, however, a theory is not asymptotically safe, then […] renormalization as described here will not be possible’ (ibid., p.275-6). [^20] Here it is worth mentioning the work of Franklin (2020) on the effectiveness of Effective Field Theories. Commenting on the paper of Butterfield (2014), Franklin notes that the appeal to the fixed point structure of the RG method in order to justify renormalization is ill suited. His claim is that RG is best seen as a mathematical framework that codifies rather than explains the renormalizability of theories in QFT, in that it allows us to mathematically establish whether or not a theory is renormalizable, but nonetheless, it does not provide a sufficient explanation of the physical reasons behind renormalization. Our focus here is on mathematical relationships rather than explanation and in Franklin’s terms, our point could be re-expressed as a statement that without a non-trivial UV fixed point the RG method cannot not successfully achieve its codifying role.
@@
Unfortunately, this rather attractive theoretical diagnostic picture is very difficult to fully implement in practice. For realistic QFTs there is no fully reliable way to explicitly calculate the beta functions to all orders. Take the example of QED. To render the introduced idealizations controllable, one needs to show that the theory is asymptotically safe by calculating the corresponding beta functions in realistic QED models, however, no rigorous proof for the asymptotic safety of QED has been achieved. Rather, what we have available are further mathematical idealization techniques for approximating the beta functions in some cases, which further suggest a possible behaviour of the theory in higher energies. Interestingly, the situation is then that we are attempting to justify the assumptions introduced by regularization and renormalization, by introducing further assumptions, creating a form of regress in which idealizations are justified by further idealizations. [^21] This is not to say that this is a vicious regress, however. Whether the justification of an approximation by means of a further approximation is epistemically warranted is an interesting open question and deserves to be studied on its own merit in future work. As already noted, the possibility for precisely this situation was anticipated in (Hartmann 2001, p. 293).
@@
In any case, even ignoring the regress worry, the evidence gained from approximate methods for calculating the beta functions is not in fact promising for the case of QED. In particular, calculations at the one-loop level indicate the existence of a Landau pole (Hollowood 2013, p.37-41) and thus provide evidence against the asymptotic safety of QED. One might hope, at this point, that the Landau pole is an artefact of the perturbation calculation of the beta function and there is some evidence in this regard (Goeckeler et al. 1998). However, no indication of a non-trivial UV fixed point for QED has been found so far. Furthermore, evidence from non-perturbative lattice simulations (Goeckeler et al. 1998; Kim et al. 2001) and the exact renormalization group approach (Gies and Jaeckel 2004) supports the triviality of the theory with complete charge screening if the UV cut-off is sent to infinity.
@@
The standard response to the presumed lack of asymptotic safety in quantum field theories found in the physics literature is that such theories should not be understood as fundamental theories and a consistent UV completion can be obtained by appeal to more fundamental theories. However, it is not expected that electroweak theory can provide a UV completion of QED and, indeed, the triviality problem is expected to also occur in the Yukawa-Higgs sector. [^22] See for example Hambye and Riesselmann (1997) A full list of references is given in Eichhorn and Held (2018) reference [4]. Rather, the hope for a UV completion of QED relies upon appeal to physics beyond the standard model, such as in asymptotically safe quantum gravity scenarios (Christiansen and Eichhorn 2017; Eichhorn 2019). Whether or not such a response to the QED triviality problem ultimately proves successful is, however, tangential to the issue at hand: we are concerned with the justification for the idealizations involved in the theories and models in the context of quantum field theories, not exogenous justifications of the idealizations coming from other theories, cf. Crowther and De Haro (2022). Moreover, since this putative justification would need to come from theories we do not in fact have, the supposed resolution has the status of a promissory note. For the case of QED the mathematical idealization in question is evidently uncontrolled (since no fixed point has been identified), and potentially also uncontrollable (since there is evidence that none in fact exists and the theory is trivial).
@@ 6. Implications
The analysis provided thus far has important implications for the philosophical understanding of theoretical models and the empirical adequacy of scientific theories. Let us recap the argument provided in Section 1. The first premise, P1, was that quantum field theory is a more general framework from which specific theories such as quantum electrodynamics and quantum chromodynamics are derived via particular action principles. The second premise, P2, was that a theory has theoretical models if its predicted empirical quantities are derivable either as deductive consequences of the theory or as a result of an approximate derivation involving controllable mathematical idealizations. The third premise, P3, was that we have good reasons to believe that the derivation of predictions from empirically successful theories in contemporary particle physics, such as quantum electrodynamics, involves ineliminable appeal to uncontrolled idealizations. These three premises together lead us to the deeply unintuitive conclusion, C1, that we have good reasons to believe that theories in contemporary particle physics, such as quantum electrodynamics, do not have theoretical models.
The natural response for those seeking to reject the unintuitive conclusion is to challenge the controversial premise P3. However, as we saw in the previous section, P3 is, as it stands, remarkably resilient, even in the face of the sophisticated framework of renormalization group theory.
An alternative response would be to appeal to the possible existence of models of QED that are related to theory via as yet unconcieved deductive relations or controlled idealizations. That is, since, logically speaking, whether theoretical models exist or not is independent of our epistemic position, such models may exist even if we are not, and indeed may never be, in a position to understand them. The plausibly of such a response will be heavily dependant on one’s background philosophy of science in particular with regard to the ontology of theories and models. If one thinks of theories and their models as real existing abstract objects then this response might allay some fears (Giere 1988; Psillos 2011). By contrast, under a quietist or pragmatic stance regarding what it is for models to exist, the response is entirely empty (French 2010; Antoniou 2021). In any case, our conclusion that contemporary particle physics includes theories which we have good reasons to believe are theories without theoretical models is unaffected by the speculative possibility that such models may exist even if we have no good reasons to believe that they do.
It might therefore seem like the only choice left is to ‘bite the bullet’ and accept the conclusion of theories without models, despite its unintuitiveness. However, as we will argue now, this option is undesirable since C1 has further clearly unacceptable consequences for empirical adequacy. Consider the following extension of our argument:
@@
    Contemporary particle physics includes theories, such as quantum electrodynamics, which we have good reasons to believe are theories without theoretical models.
@@
    A theory is empirically adequate only if it possesses theoretical models that agree with experimental outcomes.232323cf. (van Fraassen 1991, p.12): ‘To be empirically adequate is to have some model which can accommodate all the phenomena’.
@@
    We have good reasons to believe that the theory of quantum electrodynamics is not empirically adequate.
@@
Clearly, C2 is not just unintuitive but absurd and thus unacceptable. Quantum electrodynamics is variously understood as the most empirically successful theory humanity has yet constructed, so any argument that ends up with the conclusion that we have good reasons to believe it is not empirical adequate must have gone wrong somewhere! Given that our robust defence of P3 is accepted and that P1 is also accepted as uncontroversial, the two remaining options to avoid absurdity are then to put pressure on P2 and P4. Let us consider each of these options in turn.
@@
If theoretical models are not required to be connected to their background theories via controlled idealizations, then one clearly avoids the absurd conclusion since one is able to understand the empirically relevant models of quantum electrodynamics as models of the theory. The cost to the coherence of our concepts is, however, catastrophic. This is because the expansion of the definition of a theoretical model to include links via uncontrollable mathematical idealizations, simpliciter, allows arbitrary mathematical connections between scientific theories and models. To put it differently, saying that the models of a theory include models which are the products of both controllable and uncontrollable idealizations without further qualifications, allows any possible mathematical connection of any possible model with a background theory. Such a view would lead to a narrowly instrumentalist understanding of the empirical adequacy of theories which is solely defined in terms of the empirical success of the derived models, regardless of whether such models are constructed based upon rigorous and well justified scientific and mathematical processes or not. Consequently, under such view, any scientific theory in which various ad hoc mathematical modifications are introduced with the sole purpose of achieving agreement with data can be considered as empirically adequate insofar as the derived predictions agree with measurement outcomes. We do not, therefore, take the route via weakening P2 to be worth pursuing.
A plausible alternative is to distinguish between appropriate and inappropriate uncontrolled idealizations in order to avoid arbitrary mathematical simplifications of models. That is, it may be argued that the models of a theory are only those models in which the introduced mathematical idealizations, although uncontrolled, are still motivated by theoretical reasons. This approach succeeds in distinguishing between cases in which mathematical idealizations are completely arbitrary (i.e., unmotivated), explicitly known to be problematic (i.e., we have theoretical reasons against them), and cases in which the scientific community has good reasons to believe that the introduced idealizations, although uncontrolled, are plausible (i.e. motivated by by theoretical reasons). In fact, this is indeed precisely what happens in the framework of the renormalization group method where a possible justification for the truncation of the series in sought in terms of asymptotic safety and the framework of effective field theories.
This idea would lead to a more relaxed definition of empirical adequacy based on the distinction between appropriate and inappropriate uncontrolled idealizations. Such an approach would require a robust framework for distinguishing between uncontrollable idealizations that are appropriate for the derivation of empirical predictions and other arbitrary and ad hoc idealizations to make models mathematically tractable. Although intuitively some idealizations seem plainly wrong, e.g. the introduction of an arbitrary number of free parameters to fit the data, the presented analysis of the status of idealizations in quantum electrodynamics indicates that there is not as yet a clear basis to determine whether a certain uncontrolled mathematical idealization is appropriate or not in the context of QFTs. The problem is that the concept of ‘appropriate’ and ‘theoretically motivated’ idealization is extremely difficult to define in a rigorous and unambiguous way leading to a clear definition of theoretical models. What counts as appropriate and theoretically motivated depends to a large extend on the subjective appraisals of the involved scientists. We remain sympathetic to this option however, and acknowledge that further work is required to clarify these notions.
@@
In the context of the failure of Borel summability, the exogenous control approach could plausibly be understood to be illustrated by the idea of appealing to non-perturbative ‘instanton’ corrections as a means to control the idealization within the perturbative theory (Marino 2014). The key idea is to use properties of the non-perturbative solutions to understand the failure of Borel summability. In particular, we can differentiate two types of situation when there is no Borel summability: i) we are considering perturbative series around unstable minima; and ii) there are extra saddle points in the path integral. The second situation is understood to correspond to non-perturbative effects of the instanton type. In that context, one can consider ‘lateral Borel resummations’ (i.e. integrals along paths in the complex plane) and the ground state energy can be reconstructed from the Borel-resummed perturbative series corrected by Borel-resummed instanton solutions (Marino 2014, p. 37). The idea is, then, that realistic perturbative quantum field theories which are not Borel summable, admit a non-perturbative analysis with instanton corrections (or generalisations thereof). Formally, the relevant properties have only been shown to hold in toy models and thus the idealization in question is certainly not controlled as yet. However, there may be scope for experimental demonstration of the existence of instantons in the standard model (Amoroso et al. 2021; Tasevsky et al. 2023) and this would go a long way towards provision of the kind of control that we take the idealizations of perturbative quantum field theory to require in order to anchor their empirical adequacy. Furthermore, one might plausibly argue that instantons are not really exogenous: even though they amount to novel particle content and involve going beyond the perturbative framework, instantons are still part of the framework of QFT in general terms. As such, one might argue they constitute a means for endogenous control. This issue is also worthy of further analysis.
@@
There are also, as already mentioned, plausible possibilities for exogenous arguments that could provide control of the idealizations in the context of renormalization. These also come from a non-perturbative analysis but are exogenous to an even greater degree since they require us to include additional effects due to gravity. In particular, in the context of failure of asymptotic safety in quantum electrodynamics, scientists in the field of quantum gravity propose to control of the relevant idealization by appeal to the existence of a fixed point in a unified theory of quantum gravity. This wider definition would be aligned to the general attitude that UV divergences of QFT point to quantum gravity as discussed by Crowther and De Haro (2022) and to the particular motivations for the asymptotic safety programme (Eichhorn 2019). [^24] See also Crowther and Linnemann (2019).